OpenStudy (anonymous):

if the two legs have lenght 10 ins how long is the hypotenus

OpenStudy (imstuck):

Are you familiar with Pythagorean's Theorem?

OpenStudy (mathmate):

Also check if the given triangle is a right-triangle, i.e. the included angle is indeed 90 degrees

OpenStudy (imstuck):

Well if it has a hypotenuse, it is a right triangle.

OpenStudy (marissalovescats):

Welcome to OpenStudy @paeng!

OpenStudy (ipwnbunnies):

Doesn't 'hypotenuse' imply it's a right triangle?

OpenStudy (mathmate):

Thank you for your comments, and you all have a point. However, it's never too cautious to check when you have seen the number of students actually apply Pythagoras Theorem to arbitrary triangles.

OpenStudy (marissalovescats):

I'd say if the question is asking for the hypotenuse, then it's a right triangle. Especially if he doesn't have a picture, and just a statement.

OpenStudy (mathmale):

IMStuck ("Well if it has a hypotenuse, it is a right triangle") and marissalovescats (above) are correct. Out of curiosity and desire to know, I looked up "hypotenuse" on the 'Net and found that that word does indeed apply only to RIGHT triangles. When in doubt, look up the word or concept, please.

OpenStudy (mathmale):

So, the triangle at hand is an isosceles right triangle, whose shorter legs both have length 10 inches. Find the length of the hypotenuse, please.

OpenStudy (anonymous):

Use Pythagoras bro....P^2 + B^2=H^2.... Sorry Moderator can I solve this ???

OpenStudy (mathmate):

@mathmale The meaning of the word has never been disputed. What was being discussed was why should it be questioned, and I explained why.

OpenStudy (mathmale):

Or, in other words, it's imperative that would-be problem solvers determine whether or not the triangles they're trying to solve are right triangles. I'm with you; no conflict there.

OpenStudy (anonymous):

@mathmale Can I just solve this thing.... We are doing all sorts of discussions and not just helping Paeng... here

OpenStudy (mathmate):

@mathmale thank you! :)

OpenStudy (mathmate):

@paeng are you there?

OpenStudy (mathmate):

@paeng Sorry about the lengthy problem not directly related to your problem. If you have read between the lines, the suggestions to solve the problem are the following: 1. If the triangle is a right angled triangle, you can use Pythagoras Theorem \[x^2+y^2=z^2\] where z is the hypotenuse, and x, y are the lengths of the legs. Would you be able to solve for the hypotenuse?